# HyperAdaLoRA: Accelerating LoRA Rank Allocation During Training via Hypernetworks without Sacrificing Performance

### v1

@@ Abstract
Hao Zhang1thanks: These authors contribute equally to this work., Zhenjia Li211footnotemark: 1, Runfeng Bao311footnotemark: 1, Yifan Gao4, Xi Xiao5, Bo Huang2, Yuhang Wu6, Tianyang Wang5, Hao Xu1thanks: Corresponding author. 1Harvard University, 2University of Chinese Academy of Sciences 3Fudan University, 4University of Chicago, 5University of Alabama at Birmingham 6Shanghai University of Engineering Science
@@ 1 Introduction
Parameter-efficient fine-tuning (PEFT) has emerged as a practical solution for adapting large language models (LLMs) to downstream tasks by updating a small subset of parameters, thereby reducing computational and memory overhead Li and Liang (2021); Lester et al. (2021); Zaken et al. (2022); Zhang et al. (2025a, b); Houlsby et al. (2019). A prominent PEFT method, Low-Rank Adaptation (LoRA) Hu et al. (2022), is particularly notable for introducing trainable low-rank matrices into pre-trained weights during fine-tuning, reparameterizing weight updates as:
@@
where $W^{(0)},\Delta W\in\mathbb{R}^{d_{1}\times d_{2}}$, $A\in\mathbb{R}^{r\times d_{2}}$ and $B\in\mathbb{R}^{d_{1}\times r}$ with $r\ll\{d_{1},d_{2}\}$. During fine-tuning, only matrices $B$ and $A$ are updated, substantially reducing the number of trainable parameters. However, LoRA assigns a uniform rank across all layers, neglecting the varying functional importance of different components, potentially limiting performance in deep or heterogeneous models Hu et al. (2023); Zhang et al. (2023a).
@@
Dynamic rank allocation methods have been introduced to tackle this challenge by adaptively assign different ranks $r$ to various modules or layers. There are three main strategies: 1) Singular value decomposition (SVD) methods Zhang et al. (2023b); Hu et al. (2023); Zhang et al. (2023a) decompose matrices into singular values and vectors, effectively capturing key components. However, the decomposition process is computationally intensive, with a time complexity of $O(n^{3})$, and requires additional memory to store the singular values and vectors. 2) Single-rank decomposition (SRD) methods Mao et al. (2024); Zhang et al. (2024); Liu et al. (2024b) decompose matrices into rank-1 components, enabling more granular rank allocation. Despite this flexibility, identifying and pruning rank-1 components necessitates multi-stage training, increasing algorithmic complexity. The iterative selection process can also introduce instability, particularly when essential components are mistakenly pruned. 3) Rank sampling methods Valipour et al. (2022) dynamically allocates ranks during training by sampling from a range of ranks, offering post-training flexibility. However, the stochastic nature of this approach introduces gradient noise, potentially destabilizing convergence.
@@
Hypernetworks Ha et al. (2016) is a meta-model that generates parameters for a target network, decoupling parameter generation from model architecture. Some fine-tuning methods Ortiz-Barajas et al. (2024); Liang et al. (2022); Li et al. (2025) leverage hypernetworks to enhance performance on both single-task and multi-task settings. By producing task-specific parameters based on contextual inputs, hypernetworks enable dynamic adaptation without explicit iterative optimization, capturing intricate parameter adaptation patterns in real-time. Hypernetworks facilitate adaptive parameter optimization by modulating both the update direction and magnitude in response to the current model state Lorraine and Duvenaud (2018). This dynamic modulation fosters efficient navigation through high-dimensional parameter spaces, accelerating convergence while maintaining model expressiveness Kirsch et al. (2018); Shi et al. (2022).
@@
In this paper, we introduce HyperAdaLoRA, a novel framework that leverages hypernetworks Ha et al. (2016) to achieve a significant improvement in convergence speed through parameter generation. Unlike traditional methods that directly train the incremental matrices $P$, $\Lambda$, and $Q$, HyperAdaLoRA employs task-specific hypernetworks to generate these matrices. Architecturally, our hypernetworks are based on a specific attention layer of BERT Devlin et al. (2019), which enables them to capture the complex dependencies among parameters. Specifically, each hypernetwork takes the current state of $P$, $\Lambda$, or $Q$ as input and outputs their updated versions. Dynamic rank allocation is realized by pruning the output of the hypernetwork that generates $\Lambda$. The training objective of HyperAdaLoRA is to minimize the discrepancy between the parameters generated by the hypernetworks and the ideal parameters. Extensive experiments demonstrate that our method achieves faster convergence while maintaining accuracy. Additionally, further extension experiments on other LoRA-based (SRD and rank-sampling) approaches validate the broad applicability of our method.
@@ 2 Related Work
2.1 SVD-based Dynamic Rank Allocation
@@
where $P\in\mathbb{R}^{d_{1}\times r}$ and $Q\in\mathbb{R}^{r\times d_{2}}$ represent the left/right singular vectors, and the diagonal matrix $\Lambda\in\mathbb{R}^{r\times r}$ contains the singular values $\{\lambda_{i}\}_{1\leq i\leq r}$ with $r\ll\min(d_{1},d_{2})$. For example, AdaLoRA Zhang et al. (2023b) prunes less important singular values based on a sensitivity-derived importance score during training. SaLoRA Hu et al. (2023) adaptively adjusts the rank by identifying and suppressing less informative singular components, optimizing parameter efficiency across layers. IncreLoRA Zhang et al. (2023a) incrementally increases the rank during training, starting with a minimal rank and expanding as needed, balancing early training stability with later-stage expressiveness. These methods effectively capture principal components but incur $O(n^{3})$ complexity and substantial memory overhead, impacting scalability and training stability, particularly with dynamic rank adjustment.
2.2 SRD-based Dynamic Rank Allocation
SRD-based methods decompose the LoRA update $\Delta W=\sum_{i=1}^{r}u_{i}v_{i}^{T}$ into a series of rank-1 matrices, where each component $u_{i}v_{i}^{T}$ represents a distinct direction in the parameter space. This decomposition allows the model to assess and adjust each rank-1 update independently. AutoLoRA Zhang et al. (2024) uses a meta-learning scheme to determine which rank-1 slices to retain or prune, while ALoRA Liu et al. (2024b) trains a ’super-network’ and reallocates ranks based on importance. SoRA Ding et al. (2023) applies sparsity penalties to zero out less impactful components, and DoRA Liu et al. (2024a) adjusts only the direction component, assigning a learnable scalar for each weight. However, SRD methods face limitations such as increased algorithmic complexity from multi-stage training and additional optimization for rank-1 component selection. Abrupt pruning can destabilize training, while iterative selection adds computational overhead and heightens sensitivity to hyperparameter tuning.
2.3 Rank Sampling-based Dynamic Rank Allocation
Rank-sampling methods treat the LoRA rank as a random variable during training. DyLoRA Valipour et al. (2022) implements this by sampling a truncation level $b\leq R$ in each iteration, zeroing out the bottom RbR-b components and enabling the model to operate across multiple ranks without retraining. This approach eliminates the need for exhaustive rank search while also serving as a regularizer, concentrating key features in top components to potentially enhance generalization. However, training across multiple ranks can dilute performance at specific ranks compared to fixed-rank LoRA, and dynamic masking introduces slight computational overhead, potentially requiring additional training epochs to converge.
@@ 3.1 Preliminary
Hypernetworks Ha et al. (2016) are a type of neural network architecture used to generate the weights of another neural network (the target network). This can be expressed using the following formula:
@@
The input to the hypernetwork Ha et al. (2016) can be any information related to the target network, such as the input data of the target network, task requirements, or other contextual information. By conditioning parameter generation on these inputs, the hypernetwork can dynamically generate different parameters to adapt to different tasks or environments. This approach mitigates the need for learning a full set of parameters, thereby reducing model complexity and promoting generalization.
In neural architecture search (NAS), hypernetworks Ha et al. (2016) can generate parameters for multiple sub-networks, thereby efficiently exploring different network architectures. By simultaneously training the hypernetwork and the sub-networks, the performance of a large number of candidate architectures can be evaluated in a relatively short time. This allows for the rapid screening of network structures with better convergence performance. This efficient architecture exploration method helps to find more optimal model architectures, thereby accelerating the overall training and convergence process. We present a more comprehensive theoretical analysis that elucidates how hypernetworks accelerate convergence (Appendix A).
@@ 3.2 Hypernetworks Accelerate Convergence
where $\gamma>0$ denotes the regularization coefficient. This formulation ensures that the generated matrices $P$ and $Q$ approximate orthogonal transformations while maintaining compatibility with downstream tasks. We additionally provide details on the parameter matrix update process (Appendix B) and the pruning strategy (Appendix C).
@@ 3.3 Attention Driven Parameter Interaction
4 Experiments
4.1 Experimental Setup
4.1.1 Benchmarks
We conduct a comprehensive evaluation of our method, covering a wide range of tasks in both natural language understanding (NLU) and natural language generation (NLG). In the realm of natural language understanding, our method is tested on challenging tasks from the GLUE benchmark Wang (2018): RTE Wang (2018) and WNLI Wang (2018). These tasks represent large scale entailment classification, small-scale binary entailment classification, and coreference resolution presented in the form of binary entailment classification, respectively. In the natural language generation domain, we assess our method using three widely recognized datasets: Stanford Alpaca Taori et al. (2023), Magpie-Pro-300K-Filtered Xu et al. (2024), and OpenPlatypus Lee et al. (2023). In the following text, we sometimes abbreviate Magpie-Pro-300K-Filtered as Magpie.To further demonstrate the generalizability of our approach, we conduct additional evaluations on the GSM8K Cobbe et al. (2021) and HumanEval Chen et al. (2021) benchmarks.
4.1.2 Models
We use two prominent pretrained language models for NLU: RoBERTa-base Liu (2019), which is renowned for its strong performance across a wide range of NLU tasks, and DeBERTa-v3-base He et al. (2021), an enhanced version that incorporates advanced pretraining techniques. For NLG, we employ two models: LLaMA3.1-8B Grattafiori et al. (2024), a powerful 8 billion parameter model optimized for high-quality text generation, and Qwen2.5-7B Yang et al. (2024), a model that demonstrates exceptional performance in various NLG tasks. We further evaluate our method on the larger Qwen2.5-14B Yang et al. (2024).
4.1.3 Baselines
Our primary baseline for comparison is AdaLoRA Zhang et al. (2023b). AdaLoRA uses $P\Lambda Q$ as trainable parameters that are dynamically updated during training. It allocates parameter budgets by parameterizing updates in an SVD form and prunes singular values based on importance scores during training. To further demonstrate the broad applicability of our method, we additionally conduct experiments by integrating it with LoRA Hu et al. (2022), DoRA (SRD-based) Liu et al. (2024a), and DyLoRA (rank-sampling based) Valipour et al. (2022).
4.1.4 Implementation Details
Our experiments are conducted using the PyTorch framework Paszke et al. (2019) and the Hugging Face Transformers library Wolf (2020), running on a cluster equipped with NVIDIA A100 40GB GPUs. In our experiments, we set the rank $r$ to 3 and the orthogonality regularization coefficient $\gamma$ to 0.1. We use the Adam optimizer Kingma and Ba (2014) with a learning rate of $1\times 10^{-5}$ and a batch size of 64 for training. For the comparative experiments, we keep the hyperparameters for the base model fine-tuning consistent across different methods. We use ROUGE-1 and BLEU-4 as the NLG evaluation metrics. The hypernetwork is implemented as a single layer of TinyBERT, featuring a hidden dimension of 312 and 12 attention heads, with approximately 11.88 million parameters. For the CNN baseline, we adopt the ResNet-18 architecture, comprising 18 layers with hidden dimensions of 64, 128, and 256, totaling around 11.7 million parameters. Additionally, the MLP baseline consists of 15 layers with hidden dimensions of 256, 512, and 1024, amounting to approximately 11.4 million parameters. More training details can be found in Appendix D.
4.2 NLG Task Results
4.2.1 Performance Comparison
@@ 4.4.1 Convergence Time on Other LoRA-Based Methods
To demonstrate this generality, we extend it beyond AdaLoRA to several representative methods, including LoRA, DoRA (SRD-based) and DyLoRA (rank-sampling based). We evaluate these variants on LLaMA3.1-8B with a batch size of 8 using the Stanford Alpaca dataset. As shown in Table 2, our method consistently accelerates convergence across all settings.
@@ 4.4.2 Performance on Reasoning Benchmarks
Dataset LoRA AdaLoRA DoRA DyLoRA
GSM8K 93.85 93.87 94.10 94.30 94.18 94.20 94.28 94.29
HumanEval 97.00 97.03 97.63 97.64 97.56 97.60 97.39 97.42

### v2

@@ Abstract
Hao Zhang1thanks: These authors contribute equally to this work., Zhenjia Li211footnotemark: 1, Runfeng Bao311footnotemark: 1, Yifan Gao4, Xi Xiao5, Heng Zhang6, Shuyang Zhang7, Bo Huang2, Yuhang Wu8, Tianyang Wang5, Hao Xu1thanks: Corresponding author. 1Harvard University, 2University of Chinese Academy of Sciences, 3Fudan University 4University of Chicago, 5University of Alabama at Birmingham, 6South China Normal University 7Shanghai Artificial Intelligence Laboratory, 8Shanghai University of Engineering Science
@@ 1 Introduction
Parameter-efficient fine-tuning (PEFT) has emerged as a practical solution for adapting large language models (LLMs) to downstream tasks by updating a small subset of parameters, thereby reducing computational and memory overhead li2021; lester2021; zaken2022; zhang2025pdtrim; zhang2025trimtokenator; houlsby2019. A prominent PEFT method, Low-Rank Adaptation (LoRA) hu2022lora, is particularly notable for introducing trainable low-rank matrices into pre-trained weights during fine-tuning, reparameterizing weight updates as:
@@
where $W^{(0)},\Delta W\in\mathbb{R}^{d_{1}\times d_{2}}$, $A\in\mathbb{R}^{r\times d_{2}}$ and $B\in\mathbb{R}^{d_{1}\times r}$ with $r\ll\{d_{1},d_{2}\}$. During fine-tuning, only matrices $B$ and $A$ are updated, substantially reducing the number of trainable parameters. However, LoRA assigns a uniform rank across all layers, neglecting the varying functional importance of different components, potentially limiting performance in deep or heterogeneous models hu2023structure; zhang2023increlora.
@@
Dynamic rank allocation methods have been introduced to tackle this challenge by adaptively assign different ranks $r$ to various modules or layers. There are three main strategies: 1) Singular value decomposition (SVD) methods zhang2023adalora; hu2023structure; zhang2023increlora decompose matrices into singular values and vectors, effectively capturing key components. However, the decomposition process is computationally intensive, with a time complexity of $O(n^{3})$, and requires additional memory to store the singular values and vectors. 2) Single-rank decomposition (SRD) methods mao2024dora; zhang2024autolora; liu2024alora decompose matrices into rank-1 components, enabling more granular rank allocation. Despite this flexibility, identifying and pruning rank-1 components necessitates multi-stage training, increasing algorithmic complexity. The iterative selection process can also introduce instability, particularly when essential components are mistakenly pruned. 3) Rank sampling methods valipour2022dylora dynamically allocates ranks during training by sampling from a range of ranks, offering post-training flexibility. However, the stochastic nature of this approach introduces gradient noise, potentially destabilizing convergence.
@@
Hypernetworks ha2016hypernetworks is a meta-model that generates parameters for a target network, decoupling parameter generation from model architecture. Some fine-tuning methods ortiz2024hyperloader; lianghart; li2025hyperlora leverage hypernetworks to enhance performance on both single-task and multi-task settings. By producing task-specific parameters based on contextual inputs, hypernetworks enable dynamic adaptation without explicit iterative optimization, capturing intricate parameter adaptation patterns in real-time. Hypernetworks facilitate adaptive parameter optimization by modulating both the update direction and magnitude in response to the current model state lorraine2018stochastic. This dynamic modulation fosters efficient navigation through high-dimensional parameter spaces, accelerating convergence while maintaining model expressiveness kirsch2018modular; shi2022layerconnect.
@@
In this paper, we introduce HyperAdaLoRA, a novel framework that leverages hypernetworks ha2016hypernetworks to achieve a significant improvement in convergence speed through parameter generation. Unlike traditional methods that directly train the incremental matrices $P$, $\Lambda$, and $Q$, HyperAdaLoRA employs task-specific hypernetworks to generate these matrices. Architecturally, our hypernetworks are based on a specific attention layer of BERT devlin2019bert, which enables them to capture the complex dependencies among parameters. Specifically, each hypernetwork takes the current state of $P$, $\Lambda$, or $Q$ as input and outputs their updated versions. Dynamic rank allocation is realized by pruning the output of the hypernetwork that generates $\Lambda$. The training objective of HyperAdaLoRA is to minimize the discrepancy between the parameters generated by the hypernetworks and the ideal parameters. Extensive experiments demonstrate that our method achieves faster convergence while maintaining accuracy. Additionally, further extension experiments on other LoRA-based (SRD and rank-sampling) approaches validate the broad applicability of our method.
@@ 2 Related Work
2.1 SVD-based Fine-tuning Method
@@
where $P\in\mathbb{R}^{d_{1}\times r}$ and $Q\in\mathbb{R}^{r\times d_{2}}$ represent the left/right singular vectors, and the diagonal matrix $\Lambda\in\mathbb{R}^{r\times r}$ contains the singular values $\{\lambda_{i}\}_{1\leq i\leq r}$ with $r\ll\min(d_{1},d_{2})$. For example, AdaLoRA zhang2023adalora prunes less important singular values based on a sensitivity-derived importance score during training. SaLoRA hu2023structure adaptively adjusts the rank by identifying and suppressing less informative singular components, optimizing parameter efficiency across layers. IncreLoRA zhang2023increlora incrementally increases the rank during training, starting with a minimal rank and expanding as needed, balancing early training stability with later-stage expressiveness. These methods effectively capture principal components but incur $O(n^{3})$ complexity and substantial memory overhead, impacting scalability and training stability, particularly with dynamic rank adjustment.
2.2 SRD-based Fine-tuning Method
SRD-based methods decompose the LoRA update $\Delta W=\sum_{i=1}^{r}u_{i}v_{i}^{T}$ into a series of rank-1 matrices, where each component $u_{i}v_{i}^{T}$ represents a distinct direction in the parameter space. This decomposition allows the model to assess and adjust each rank-1 update independently. AutoLoRA zhang2024autolora uses a meta-learning scheme to determine which rank-1 slices to retain or prune, while ALoRA liu2024alora trains a ’super-network’ and reallocates ranks based on importance. SoRA ding2023sparse applies sparsity penalties to zero out less impactful components, and DoRA liu2024dora adjusts only the direction component, assigning a learnable scalar for each weight. However, SRD methods face limitations such as increased algorithmic complexity from multi-stage training and additional optimization for rank-1 component selection. Abrupt pruning can destabilize training, while iterative selection adds computational overhead and heightens sensitivity to hyperparameter tuning.
2.3 Rank Sampling-based Fine-tuning Method
Rank-sampling methods treat the LoRA rank as a random variable during training. DyLoRA valipour2022dylora implements this by sampling a truncation level $b\leq R$ in each iteration, zeroing out the bottom RbR-b components and enabling the model to operate across multiple ranks without retraining. This approach eliminates the need for exhaustive rank search while also serving as a regularizer, concentrating key features in top components to potentially enhance generalization. However, training across multiple ranks can dilute performance at specific ranks compared to fixed-rank LoRA, and dynamic masking introduces slight computational overhead, potentially requiring additional training epochs to converge. Additionally, we note that some works li2025beyond; li2025beyond2 also introduce the concepts of non-zero initialization dynamics and token-wise input–output projections.
@@ 3.1 Preliminary
Hypernetworks ha2016hypernetworks are a type of neural network architecture used to generate the weights of another neural network (the target network). This can be expressed using the following formula:
@@
The input to the hypernetwork ha2016hypernetworks can be any information related to the target network, such as the input data of the target network, task requirements, or other contextual information. By conditioning parameter generation on these inputs, the hypernetwork can dynamically generate different parameters to adapt to different tasks or environments. This approach mitigates the need for learning a full set of parameters, thereby reducing model complexity and promoting generalization.
In neural architecture search (NAS), hypernetworks ha2016hypernetworks can generate parameters for multiple sub-networks, thereby efficiently exploring different network architectures. By simultaneously training the hypernetwork and the sub-networks, the performance of a large number of candidate architectures can be evaluated in a relatively short time. This allows for the rapid screening of network structures with better convergence performance. This efficient architecture exploration method helps to find more optimal model architectures, thereby accelerating the overall training and convergence process. We present a more comprehensive theoretical analysis that elucidates how hypernetworks accelerate convergence (Appendix A).
@@ 3.2 Hypernetworks Accelerate Convergence
where $\gamma>0$ denotes the regularization coefficient. This formulation ensures that the generated matrices $P$ and $Q$ approximate orthogonal transformations while maintaining compatibility with downstream tasks. We additionally provide details on the parameter matrix update process (Appendix C) and the pruning strategy (Appendix D).
@@ 3.3 Attention Driven Parameter Interaction
4 Experiments
4.1 Experimental Setup
4.1.1 Benchmarks
We conduct a comprehensive evaluation of our method, covering a wide range of tasks in both natural language understanding (NLU) and natural language generation (NLG). In the realm of natural language understanding, our method is tested on challenging tasks from the GLUE benchmark wang2018glue: RTE wang2018glue and WNLI wang2018glue. These tasks represent large scale entailment classification, small-scale binary entailment classification, and coreference resolution presented in the form of binary entailment classification, respectively. In the natural language generation domain, we assess our method using three widely recognized datasets: Stanford Alpaca alpaca, Magpie-Pro-300K-Filtered xu2024magpie, and OpenPlatypus platypus2023. In the following text, we sometimes abbreviate Magpie-Pro-300K-Filtered as Magpie.To further demonstrate the generalizability of our approach, we conduct additional evaluations on the GSM8K cobbe2021training and HumanEval chen2021evaluating benchmarks.
4.1.2 Models
We use two prominent pretrained language models for NLU: RoBERTa-base liu2019roberta, which is renowned for its strong performance across a wide range of NLU tasks, and DeBERTa-v3-base he2021debertav3, an enhanced version that incorporates advanced pretraining techniques. For NLG, we employ two models: LLaMA3.1-8B grattafiori2024llama, a powerful 8 billion parameter model optimized for high-quality text generation, and Qwen2.5-7B yang2024qwen2, a model that demonstrates exceptional performance in various NLG tasks. We further evaluate our method on the larger Qwen2.5-14B yang2024qwen2.
4.1.3 Baselines
Our primary baseline for comparison is AdaLoRA zhang2023adalora. AdaLoRA uses $P\Lambda Q$ as trainable parameters that are dynamically updated during training. It allocates parameter budgets by parameterizing updates in an SVD form and prunes singular values based on importance scores during training. To further demonstrate the broad applicability of our method, we additionally conduct experiments by integrating it with LoRA hu2022lora, DoRA (SRD-based) liu2024dora, and DyLoRA (rank-sampling based) valipour2022dylora.
4.1.4 Implementation Details
Our experiments are conducted using the PyTorch framework paszke2019pytorch and the Hugging Face Transformers library wolf2020transformers, running on a cluster equipped with NVIDIA A100 40GB GPUs. In our experiments, we set the rank $r$ to 3 and the orthogonality regularization coefficient $\gamma$ to 0.1. We use the Adam optimizer kingma2014adam with a learning rate of $1\times 10^{-5}$ and a batch size of 64 for training. For the comparative experiments, we keep the hyperparameters for the base model fine-tuning consistent across different methods. We use ROUGE-1 and BLEU-4 as the NLG evaluation metrics. The hypernetwork is implemented as a single layer of TinyBERT, featuring a hidden dimension of 312 and 12 attention heads, with approximately 11.88 million parameters. For the CNN baseline, we adopt the ResNet-18 architecture, comprising 18 layers with hidden dimensions of 64, 128, and 256, totaling around 11.7 million parameters. Additionally, the MLP baseline consists of 15 layers with hidden dimensions of 256, 512, and 1024, amounting to approximately 11.4 million parameters. More training details can be found in Appendix E.
4.2 NLG Task Results
4.2.1 Performance Comparison
@@ 4.4.1 Convergence Time on Other LoRA-Based Methods
To demonstrate this generality, we extend it beyond AdaLoRA to several representative methods, including LoRA, DoRA and DyLoRA. We evaluate these variants on LLaMA3.1-8B with a batch size of 8 using the Stanford Alpaca dataset. As shown in Table 2, our method consistently accelerates convergence across all settings. We provide the details of hypernetworks for LoRA, DoRA, and DyLoRA in Appendix B. Additional results are provided in Appendix G.
@@
Dataset LoRA AdaLoRA DoRA DyLoRA
GSM8K 71.19 71.21 72.38 72.54 72.45 72.46 72.52 72.53
HumanEval 42.26 42.27 43.54 43.56 43.51 43.53 43.43 43.45